OpenStudy (anonymous):

c+(4-3c)-2=0 help me solve this plez

OpenStudy (anonymous):

Okay so use BIDMAS Brackets Indices (powers) Division Multiplication Addition and finally then Subtraction i.e. first expand the bracket and take it from there

OpenStudy (anonymous):

\[c+(4-3c)-2=0 => c+4-3c-2=0 => -2c+2 => 2c=2 => c=1\]

OpenStudy (anonymous):

-2c+2=0 is what i meant to write on the third to last implication